# Audio Flamingo: A Novel Audio Language Model with Few-Shot Learning and Dialogue Abilities

### v2

@@ Abstract
Augmenting large language models (LLMs) to understand audio – including non-speech sounds and non-verbal speech – is critically important for diverse real-world applications of LLMs. In this paper, we propose Audio Flamingo, a novel audio language model with 1) strong audio understanding abilities, 2) the ability to quickly adapt to unseen tasks via in-context learning and retrieval, and 3) strong multi-turn dialogue abilities. We introduce a series of training techniques, architecture design, and data strategies to enhance our model with these abilities. Extensive evaluations across various audio understanding tasks confirm the efficacy of our method, setting new state-of-the-art benchmarks. ${}^{\dagger}$. Correspondence to: Zhifeng Kong (zkong@nvidia.com), Wei Ping (wping@nvidia.com), Rafael Valle (rafaelvalle@nvidia.com).
Copyright 2024 by the author(s).
@@ 1 Introduction
Our contribution to expand LLM’s ability to understand sound is called Audio Flamingo: a novel audio language model that supports in-context learning (ICL), retrieval augmented generation (RAG), and multi-turn dialogues. It achieves state-of-the-art results on multiple audio understanding tasks.  [^1] We will open-source the training and inference code. Our demo website is: https://audioflamingo.github.io/.
@@ 2 Related work
Multimodal LLMs. There has been tremendous progress in the area of multimodal LLMs. In addition to text, these models take inputs from various modalities such as vision (Tsimpoukelli et al., 2021; Alayrac et al., 2022; Yang et al., 2023; Driess et al., 2023; Liu et al., 2023a; Li et al., 2023a), audio (Deshmukh et al., 2023; Gong et al., 2023b; Rubenstein et al., 2023), or multiple of them (Han et al., 2023; Tang et al., 2023b; Moon et al., 2023; Zhao et al., 2023), and each has a different integration method. In the audio modality, prior works have looked at speech tasks (Chen et al., 2023; Rubenstein et al., 2023), general audio understanding (Deshmukh et al., 2023; Gong et al., 2023c), music understanding (Gardner et al., 2023; Won et al., 2023; Li et al., 2023c; Liu et al., 2023b; Doh et al., 2023), or a combination of these (Gong et al., 2023b; Tang et al., 2023a; Chu et al., 2023). The focus of our paper is audio understanding, which includes non-speech sound and music, and non-verbal speech. Different from prior works, our model has stronger audio understanding ability, and is the first audio understanding model with i) in-context few-shot learning ability, ii) retrieval augmented generation ability, and iii) strong multi-turn dialogue ability.
Audio encoders and representation. Many audio encoders extract audio features from the spectrogram, including CNN-based method (Kong et al., 2020) and Transformer-based methods (Gong et al., 2021; Chen et al., 2022; Défossez et al., 2022; Radford et al., 2023; Gong et al., 2023a). These methods are primarily targeted at solving a particular problem such as speech recognition or event detection. Based on these encoders, many joint audio-language embeddings have been proposed (Elizalde et al., 2023a, b; Wu et al., 2023; Huang et al., 2022; Li et al., 2023c). These methods use contrastive learning to map audio and language embeddings into the same space, and are often trained on a large variety of audio and language. However, many of these methods compute a single embedding for an audio and therefore may lose temporal information. In this paper, we build an audio encoder with sliding windows based on ClapCap (Elizalde et al., 2023b) to better capture long-range and temporal information.
@@ 3.1 Architecture
i) Audio feature extractor with sliding window. We use ClapCap (Elizalde et al., 2023b) as the audio feature extractor backbone, which we denote as $\mathcal{E}$. ClapCap takes 7-second 44.1kHz raw audio as input, then transforms the audio into Mel-spectrogram of hop length 320, window length 1024, 64 Mel bins, and finally outputs a 1024-dimensional vector representation.
We consider each 7-second segment as a window and use sliding windows to extract features for longer audio. The overlap between consecutive windows is $7\times 0.75=5.25$ seconds. We use a maximum of $m=16$ sliding windows. Formally, let $x(s\texttt{:}t)$ be the segment of $s$ to $t$ seconds in audio $x$. Then, the extracted feature is $\left[\mathcal{E}(x(0\texttt{:}7)),\mathcal{E}(x(\frac{7}{4}\texttt{:}\frac{7\times 5}{4})),\cdots,\mathcal{E}(x(\frac{7(m-1)}{4}\texttt{:}\frac{7(m+3)}{4})\right]$, which supports a maximum of $33.25$ seconds. The goal of this design is to capture long-range and temporal information that might be ignored in a single fused representation vector (Wu et al., 2023). Long audio will be cropped and short audio will be zero-padded. If an entire segment is zero-padded then we will mask the corresponding embedding at cross attention. If the input is interleaved data with $>1$ audio, we concatenate their sliding window representations.
@@ 3.2 Training Method
ii) Cross attention masks. We use block triangular cross attention masks for interleaved samples so that the likelihood of $j$-th output $p_{\theta}(y_{\mathrm{out}}^{j})$ is conditioned only on the first $j$ audio inputs $x^{\leq j}$. We expect this helps making the model to look at previous audio. Figure 3 demonstrates the mask.
Figure 3: Left: the block triangular cross attention masks between text tokens and audio embeddings. Right: the loss mask of an interleaved training sample.
iii) Two training stages. We divide training into pre-training and supervised fine-tuning (SFT), a widely adopted and stable method in training LMs (Ouyang et al., 2022). During pre-training we only train the audio representation transformation layers and the gated xattn-dense layers. The purpose is to obtain a good set of initialization weights for these layers. During SFT, we unfreeze the entire LM, and train all modules but the audio encoder.  [^2] In initial experiments we found unfreezing the audio encoder led to much worse results.
@@ 4.1 Datasets
Dataset sources. We train our model on a variety of audio datasets that can be roughly classified into three types: music, non-speech general sound, and non-verbal speech. In this paper, we focus on these data given the immediate availability of state-of-the-art speech transcription models. We look at three types of tasks: (1) audio captioning (CAP), where we would like the model to describe the audio in a sentence; (2) audio question-answering (AQA), where we would like the model to answer questions regarding the audio, and (3) audio classification (CLS), where we would like the model to classify the sound into one or more labels corresponding to events, scenes, music genres, instruments, qualities, and others. An overview of all training datasets is shown in Table 6 in the Appendix.
@@ 5.1 Experimental Setup
NonSpeech17k CLS ACC $79.0\%$ (Rashid et al., 2023) $\mathbf{85.1\%}$
@@ 5.2 Q1: Strong Audio Understanding Ability
We evaluate our model on several in-distribution (train-test) benchmarks, and compare to state-of-the-art audio language model baselines. The results are shown in Table 1. Note that we define F1approx to measure inexact but similar predicted labels in FSD50k, where we consider the prediction to be correct if SentenceBERT_similarity(output, ground_truth) $>0.8$ (Reimers & Gurevych, 2019, 2020). This metric is applied to outputs from baselines as well.
@@ 5.4 Q3: Multi-Turn Dialogues
We measure Audio Flamingo’s ability to answer questions in a multi-turn dialogue setting. The context is more complex and strongly correlated between rounds (e.g. there exist many pronouns and follow-up questions). We fine-tune Audio Flamingo on the two sets that we generated (AF-Dialogue-AudioSetSL and AF-Dialogue-MusicCaps) to obtain a chat model. We evaluate the chat model on the test split of these two dialogue datasets. We take user instructions and let the model generate answers turn-by-turn (where previous generated answers become the chatting history for next generation). We compare to Qwen-Audio (Chu et al., 2023) and LTU (Gong et al., 2023c) in Table 5.  [^3] While the baseline methods claimed to support multi-turn dialogues, we were unable to find quantitative evidence. Our chat model achieves significantly better results than baseline methods. In Appendix C.4, we use qualitative samples to show that our chat model captures context such as prior information and pronouns better.
@@
Method RECAP Ours Ours
# Shots $4$ $4$ $8$
CIDEr $\uparrow$ $0.362$ $0.518$ $\mathbf{0.538}$
Table 5: Evaluation of Audio Flamingo versus baseline methods on the multi-turn dialogue test sets. A stands for AF-Dialogue-AudioSetSL and M stands for AF-Dialogue-MusicCaps. We report CIDEr, Bleu4 (Papineni et al., 2002), and Rouge-L (R-L) (Lin, 2004). Audio Flamingo significantly outperforms larger baseline models, indicating strong dialogue ability of our proposed model.
@@
Figure 5: Relative improvement of few-shot results over zero-shot results under different number of ICL samples. Using more ICL samples consistently improves few-shot results, and the benefit is dataset-dependent.
@@ 5.5 Q4: Ablation Studies
Effect of number of few-shot samples. We reduce the number of in-context few-shot samples to 4 and evaluate on the few-shot benchmarks. In Figure 5, we plot the relative improvements over zero-shot results: $(\mathrm{ACC_{few\text{-}shot}}-\mathrm{ACC_{0\text{-}shot}})/\mathrm{ACC_{0\text{-}shot}}$. We find adding more ICL samples consistently improves few-shot results, and the benefit is dataset-dependent.
Effect of LM. We found using an instruction-tuned LLM to be key to the success of audio understanding. In early experiments we investigated opt-1.3b/2.7b (Zhang et al., 2022) and phi-1.5 (Li et al., 2023b) and found the audio understanding results to be slightly but consistently worse.
Effect of audio encoder. In early experiments, we tested all variants of LAION-CLAP (Wu et al., 2023) and Microsoft-CLAP (Elizalde et al., 2023a, b) as the audio encoder backbone. We found ClapCap (Elizalde et al., 2023b) performs the best overall – in terms of both evaluation results and computing speed – whereas a large LAION-CLAP trained on more music data led to better accuracy on music information retrieval.
@@ 6 Conclusion and Future Work
In this paper, we present Audio Flamingo, an audio language model with a series of innovations that achieves the state-of-the-art results on several close-ended and open-ended audio understanding tasks without task specific fine-tuning. It also has strong ICL and RAG abilities, and has the state-of-the-art few-shot learning results. Furthermore, we design a dataset generation strategy and introduce two dialogue datasets, and enable Audio Flamingo to chat with a user and achieve state-of-the-art results on dialogue benchmarks.
One important future direction is to investigate scaling strategies for using larger LMs. Assuming that larger LMs could have better knowledge and stronger ability to follow instructions, we believe that Audio Flamingo could benefit from a larger LM. A second future direction is to investigate complex speech-related tasks beyond transcription. A third future direction is to combine the audio understanding abilities of our model with visual language models such that one model could understand video with accompanied audio.
@@ Acknowledgement
We thank Siddharth Gururani, Zihan Liu, Mostofa Patwary, Shrimai Prabhumoye, and Chen Zhu for helpful discussions. We thank Ke Chen and Yuan Gong for help on sharing datasets.
@@ Appendix A Dataset Staging, Weights, and Templates
Table 6 includes an overview of datasets (by type) we use to train Audio Flamingo.
Audio Type Task Datasets #Audio-Text Pairs
SoundDescs (Oncescu et al., 2021), Clotho-v2 (Drossos et al., 2020),
WavText5K (Deshmukh et al., 2022), LAION-630k (Wu et al., 2023)
CochlScene (Jeong & Park, 2022), NonSpeech7K (Rashid et al., 2023),
AQA MusicQA (Liu et al., 2023b), MusicAVQA (Li et al., 2022) $\sim$94 K
FMA (Defferrard et al., 2016), MusDB-HQ (Rafii et al., 2019),
MELD (Poria et al., 2018), OMGEmotion (Barros et al., 2018)
We construct instructions for each task and dataset. Below are all instruction templates we use.
@@ B.2 Prompts
Below we show the prompts used to generate AF-Dialogue-AudioSetSL and AF-Dialogue-MusicCaps.

### v3

@@ Abstract
Augmenting large language models (LLMs) to understand audio – including non-speech sounds and non-verbal speech – is critically important for diverse real-world applications of LLMs. In this paper, we propose Audio Flamingo, a novel audio language model with 1) strong audio understanding abilities, 2) the ability to quickly adapt to unseen tasks via in-context learning and retrieval, and 3) strong multi-turn dialogue abilities. We introduce a series of training techniques, architecture design, and data strategies to enhance our model with these abilities. Extensive evaluations across various audio understanding tasks confirm the efficacy of our method, setting new state-of-the-art benchmarks. Our demo website is https://audioflamingo.github.io/ and the code is open-sourced at https://github.com/NVIDIA/audio-flamingo.   1NVIDIA, CA, USA. Correspondence to: Zhifeng Kong <zkong@nvidia.com>, Wei Ping <wping@nvidia.com>, Rafael Valle <rafaelvalle@nvidia.com>.
Proceedings of the $\mathit{41}^{st}$ International Conference on Machine Learning, Vienna, Austria. PMLR 235, 2024. Copyright 2024 by the author(s).
Audio Language Model, Audio Understanding, In Context Learning, Retrieval Augmented Generation, Dialogue
@@ 1 Introduction
Our contribution to expand LLM’s ability to understand sound is called Audio Flamingo: a novel audio language model that supports in-context learning (ICL), retrieval augmented generation (RAG), and multi-turn dialogues. It achieves state-of-the-art results on multiple audio understanding tasks.
@@ 2 Related work
Multimodal LLMs. There has been tremendous progress in the area of multimodal LLMs. In addition to text, these models take inputs from various modalities such as vision (Tsimpoukelli et al., 2021; Alayrac et al., 2022; Yang et al., 2023; Driess et al., 2023; Liu et al., 2023a; Li et al., 2023a), audio (Deshmukh et al., 2023; Gong et al., 2023b; Rubenstein et al., 2023), or multiple of them (Han et al., 2023; Tang et al., 2023b; Moon et al., 2023; Zhao et al., 2023), and each has a different integration method. In the audio modality, prior works have looked at speech tasks (Chen et al., 2023; Rubenstein et al., 2023), general audio understanding (Deshmukh et al., 2023; Gong et al., 2023c), music understanding (Gardner et al., 2023; Won et al., 2023; Li et al., 2023b; Liu et al., 2023b; Doh et al., 2023), or a combination of these (Gong et al., 2023b; Tang et al., 2023a; Chu et al., 2023). The focus of our paper is audio understanding, which includes non-speech sound and music, and non-verbal speech. Different from prior works, our model has stronger audio understanding ability, and is the first audio understanding model with i) in-context few-shot learning ability, ii) retrieval augmented generation ability, and iii) strong multi-turn dialogue ability.
Audio encoders and representation. Many audio encoders extract audio features from the spectrogram, including CNN-based method (Kong et al., 2020) and Transformer-based methods (Gong et al., 2021; Chen et al., 2022; Défossez et al., 2022; Radford et al., 2023; Gong et al., 2023a). These methods are primarily targeted at solving a particular problem such as speech recognition or event detection. Based on these encoders, many joint audio-language embeddings have been proposed (Elizalde et al., 2023a, b; Wu et al., 2023; Huang et al., 2022; Li et al., 2023b). These methods use contrastive learning to map audio and language embeddings into the same space, and are often trained on a large variety of audio and language. However, many of these methods compute a single embedding for an audio and therefore may lose temporal information. In this paper, we build an audio encoder with sliding windows based on ClapCap (Elizalde et al., 2023b) to better capture long-range and temporal information.
@@ 3.1 Architecture
i) Audio feature extractor with sliding window. We use ClapCap (Elizalde et al., 2023b) as the audio feature extractor backbone, which we denote as $\mathcal{E}$. ClapCap is hard-coded to take 7-second of 44.1kHz raw audio as input, then transforms the audio into Mel-spectrogram of hop length 320, window length 1024, 64 Mel bins, and finally outputs a 1024-dimensional vector representation.
We consider each 7-second segment as a window and use sliding windows to extract features for longer audio. The overlap between consecutive windows is $7\times 0.75=5.25$ seconds. Formally, let $x(s\texttt{:}t)$ be the segment of $s$ to $t$ seconds in audio $x$. Then, the extracted feature is $\left[\mathcal{E}(x(0\texttt{:}7)),\mathcal{E}(x(\frac{7}{4}\texttt{:}\frac{7\times 5}{4})),\cdots,\mathcal{E}(x(\frac{7(m-1)}{4}\texttt{:}\frac{7(m+3)}{4})\right]$. The intuition of this design is to capture long-range and temporal information that might be ignored in a single fused representation vector (Wu et al., 2023). We use a maximum of $m=16$ sliding windows, which supports a maximum of $33.25$ second audio length.  [^1] We use $m=16$ because most training samples are $<30$s. Note that our cross attention complexity is linear in $m$ and therefore the audio length. In prior self-attention methods the attention complexity is quadratic in the audio length. Long audio will be cropped and short audio will be zero-padded. If an entire segment is zero-padded then we will mask the corresponding embedding at cross attention. If the input is interleaved data with $>1$ audio, we concatenate their sliding window representations.
@@ 3.2 Training Method
ii) Cross attention masks. We use block upper-triangular cross attention masks for interleaved samples so that the likelihood of $j$-th output $p_{\theta}(y_{\mathrm{out}}^{j})$ is conditioned only on the first $j$ audio inputs $x^{\leq j}$. We expect this helps making the model to look at previous audio. Figure 3 demonstrates the mask.
Figure 3: Left: the block upper-triangular cross attention masks between text tokens and audio embeddings. Right: the loss mask of an interleaved training sample.
iii) Two training stages. We divide training into pre-training and supervised fine-tuning (SFT), a widely adopted and stable method in training LMs (Ouyang et al., 2022). During pre-training we only train the audio representation transformation layers and the gated xattn-dense layers. The purpose is to obtain a good set of initialization weights for these layers. During SFT, we unfreeze the entire LM, and train all modules but the audio encoder.  [^2] In initial experiments we found unfreezing the audio encoder caused training instability.
@@ 4.1 Datasets
Dataset sources. We train our model on a variety of audio datasets that can be roughly classified into three types: music, non-speech general sound, and non-verbal speech. In this paper, we focus on these data given the immediate availability of state-of-the-art speech transcription models. We look at three types of tasks: (1) audio captioning (CAP), where we would like the model to describe the audio in a sentence; (2) audio question-answering (AQA), where we would like the model to answer questions regarding the audio, and (3) audio classification (CLS), where we would like the model to classify the sound into one or more labels corresponding to events, scenes, music genres, instruments, qualities, and others. An overview of all training datasets is shown in Table 1.
Audio Type Task Datasets #Audio-Text Pairs
SoundDescs (Oncescu et al., 2021), Clotho-v2 (Drossos et al., 2020),
WavText5K (Deshmukh et al., 2022), LAION-630k (Wu et al., 2023)
CochlScene (Jeong & Park, 2022), NonSpeech7K (Rashid et al., 2023),
AQA MusicQA (Liu et al., 2023b), MusicAVQA (Li et al., 2022) $\sim$94 K
FMA (Defferrard et al., 2016), MusDB-HQ (Rafii et al., 2019),
MELD (Poria et al., 2018), OMGEmotion (Barros et al., 2018)
@@ 5.1 Experimental Setup
NonSpeech7k CLS ACC $79.0\%$ (Rashid et al., 2023) $\mathbf{85.1\%}$
@@ 5.2 Q1: Strong Audio Understanding Ability
We evaluate our model on several in-distribution (train-test) benchmarks, and compare to state-of-the-art audio language model baselines. The results are shown in Table 2. Note that we define F1approx to measure inexact but similar predicted labels in FSD50k, where we consider the prediction to be correct if the sentence BERT similarity between output and ground truth is $>0.8$ (Reimers & Gurevych, 2019, 2020). This metric is applied to outputs from baselines as well.
@@ 5.4 Q3: Multi-Turn Dialogues
We measure Audio Flamingo’s ability to answer questions in a multi-turn dialogue setting. The context is more complex and strongly correlated between rounds (e.g. there exist many pronouns and follow-up questions). We fine-tune Audio Flamingo on the two sets that we generated (AF-Dialogue-AudioSetSL and AF-Dialogue-MusicCaps) to obtain a chat model. We evaluate the chat model on the test split of these two dialogue datasets. We take user instructions and let the model generate answers turn-by-turn (where previous generated answers become the chatting history for next generation). We compare to Qwen-Audio (Chu et al., 2023), LTU (Gong et al., 2023c), and MU-LLaMA (Liu et al., 2023b) in Table 6.  [^3] While the baseline methods claimed to support multi-turn dialogues, we were unable to find quantitative evidence. Our chat model achieves significantly better results than baseline methods. In Appendix C.5, we use qualitative samples to show that our chat model captures context such as prior information and pronouns better.
@@
Dataset Ours (16-shot) $\uparrow$ $\Delta$uparrow
CREMA-D $35.4\%$ $8.9\%$
Ravdess $40.2\%$ $19.3\%$
US8K $91.9\%$ $16.9\%$
GTZAN $76.3\%$ $8.4\%$
Medley-solos-DB $96.0\%$ $3.3\%$
Method RECAP Ours Ours Ours
# Shots $4$ $4$ $8$ $16$
CIDEr $\uparrow$ $0.359$ $0.518$ $0.538$ $\mathbf{0.546}$
@@
MH LTU $0.558$ $0.083$ $0.347$
MH Ours $\mathbf{1.350}$ $\mathbf{0.207}$ $\mathbf{0.448}$
(a) Relative improvement of few-shot results over zero-shot results under different number of ICL samples. Using more ICL samples consistently improves few-shot results, and the benefit is dataset-dependent.
(b) Relative difference of results from different LM backbones and audio encoders. The opt-1.3b model without instruction-tuning is systematically worse than the instruction-tuned opt-iml-max-1.3b. LAION-CLAP is worse than ClapCap in most cases. Clap2023 is better than ClapCap in close-ended tasks, but worse in open-ended tasks including captioning and open question-answering.
(c) Relative difference of results from different training datasets compared to the Pengi baseline. Audio Flamingo achieves better results than Pengi even if no extra data is used. Larger scale pretraining and SFT lead to better results on these benchmarks on average.
@@ 5.5 Q4: Ablation Studies
Effect of number of few-shot samples. We study different numbers of in-context few-shot samples and evaluate on the few-shot benchmarks. In Figure 5(a), we plot the relative improvements over zero-shot results. Results show a clear trend that having more ICL samples improves few-shot results, and the improvement highly depends on the dataset.
Effect of architecture. In Figure 5(b), we compare results between opt-1.3b (Zhang et al., 2022) and the instruction-tuned opt-iml-max-1.3b (Iyer et al., 2022), and also compare different audio encoders including ClapCap (Elizalde et al., 2023b), Clap2023 (Elizalde et al., 2023b), and LAION-CLAP (Wu et al., 2023). In terms of the LM backbone, the instruction-tuned opt is better in most tasks. As of audio encoders, LAION-CLAP is worse in most tasks, ClapCap is better in open-ended tasks, and Clap2023 is better in most close-ended tasks.
Effect of training data. In Figure 5(c), we compare Audio Flamingo trained on three different training sets to the Pengi baseline: (1) no pretraining, SFT on our best available dataset that is strictly a subset of Pengi’s training set, (2) pretraining and SFT on this strict subset, and (3) pretraining and SFT on our curated datasets. Audio Flamingo achieves better evaluation results than Pengi even if no extra data is used. In addition, increasing the data amount in pretraining and SFT can improve the results on average.
@@ 6 Conclusion and Future Work
In this paper, we present Audio Flamingo, an audio language model with a series of innovations that achieves the state-of-the-art results on several close-ended and open-ended audio understanding tasks without task specific fine-tuning. It also has strong ICL and RAG abilities, and has the state-of-the-art few-shot learning results. Furthermore, we design a dataset generation strategy and introduce two dialogue datasets, enabling Audio Flamingo to chat with a user about the audio for multiple rounds and achieve state-of-the-art results on dialogue benchmarks.
Our model has several limitations, which we plan to address in future work. One important future direction is to investigate scaling strategies for using larger LMs. Assuming that larger LMs could have better knowledge and stronger ability to follow instructions, we believe that Audio Flamingo could benefit from a larger LM. A second future direction is to investigate complex speech-related tasks beyond transcription. This requires our model to condition on dense embeddings (Chen et al., 2023). This modification is straightforward in Audio Flamingo as the architecture is flexible enough to support new embeddings through the addition of new cross-attention heads. Another future direction is to build a audio language model that can output both text and audio and follow more complex interleaved instructions. Finally, a future direction towards unifying more modalities is to combine the audio understanding abilities of our model with visual language models (Alayrac et al., 2022) such that one model could understand image, video, and audio.
@@ Acknowledgement
We thank Siddharth Gururani, Zihan Liu, Mostofa Patwary, Shrimai Prabhumoye, and Chen Zhu for helpful discussions. We thank Ke Chen and Yuan Gong for help on sharing datasets with us.
Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning. The proposed method facilitates automation in the audio-language domain and may be used in a variety of scenarios such as education, healthcare, environment, industry, music, etc. Careful use of the model is essential to ensure compliance with copyright restrictions in certain situations.
@@ Appendix A Dataset Staging, Weights, and Templates
Table 1 includes an overview of datasets (by type) we use to train Audio Flamingo. We construct instructions for each task and dataset. Below are all instruction templates we use.
@@ B.2 Prompts
Below we show the prompts used to generate audio-based multi-turn dialogues from GPT-4 for the strongly labeled AudioSet-SL dataset (Gemmeke et al., 2017; Hershey et al., 2021) and the MusicCaps dataset (Agostinelli et al., 2023) respectively.